OpenStudy (anonymous):

Which equation represents the sentence? Eight less than a number is one third the number. (Points : 1) 8-n=n/3 n-8=3n n-8=n/3 n/3-8=n

OpenStudy (anonymous):

ok so i think the answer is c

OpenStudy (anonymous):

Eight less than a number n-8 is = one third the number. n/3

OpenStudy (anonymous):

so, your third choice is correct

OpenStudy (anonymous):

Let n be the number . Now 8 less than the number means n-8 and one third the number is n/3. Hence the required equation is n-8=n/3

OpenStudy (anonymous):

yea i think she all ready got it